All services operational.
August 3, 2022 9:45 AM | Cable Internet issue affecting customers within a region of Toronto has now been resolved. If you are still experiencing issues, please call our support line.
August 3, 2022 8:00 AM | Cable Internet customers within a region of Toronto are continuing to experience connectivity issues. Our technicians are still working on resolving the issue and will provide an ETA shortly.
August 2, 2022 8:00 PM | Cable Internet customers within a region of Toronto are experiencing issues with their Internet service. Our technicians are investigating and will provide an update when the issue is resolved.
July 11, 2022 8:00 AM EST | As per Rogers, the network outage is now resolved and Internet services have been restored.
July 8, 2022 8:00 AM EST | Services impacted by Rogers outage: Cable Internet, some Fibre Internet business customers.
July 8, 2022 8:00 AM EST | Internet services in Ontario region were impacted due to Nation-wide outage with Carrier Rogers.
Apr 21, 2022 2:30 PM EST | Carrier confirmed regional outage has been resolved.
Apr 21, 2022 11:00 AM EST | Diallog SIP service is operational, this issue has been identified as related to carrier outages at various exchanges in ON and QC. Will share updates as soon as the carrier provides them.
Apr 21, 2022 11:00 AM EST | Reports that certain customers in Ottawa region are experiencing inbound call issues.
Apr 12, 2022 12:33 PM EST | Voice services restored.
Apr 12, 2022 12:02 PM EST | We are aware that there is an issue with voice connections. Engineers are working to correct issues. Expected ETA: 30 minutes.
Apr 2, 2022 03:00 AM EST | Maintenance window to apply patches and configuration changes to bring back online all nodes and all defence mechanisms.
Apr 1, 2022 06:15 PM EST | Second DOS attack occurred, handled within 15 minutes by our engineers.
Apr 1, 2022 04:00 PM EST | Voice service restored.
Apr 1, 2022 12:00 PM EST | Intermittent performance experienced across our network with calls being handled by our failover systems while our engineers are working to neutralize the sustained DOS attacks.
Apr 1, 2022 09:00 AM EST | Issue identified as related to a DDOS attack. Failover systems overwhelmed and also suffering performance issues.
Apr 1, 2022 06:00 AM EST | Degradation of voice services experienced. Engineers investigating.
May 25, 2021 10:45 AM EST | A subset of DSL customers are experiencing authentication issues causing red Internet light on the modem. Please unplug your modem, wait 30 seconds, plug it back in, wait 3-5 minutes for green DSL and Internet light, and service should be restored.
May 24, 2021 2:00 PM EST | Service restored. Communication issue discovered between core equipment and upstream provider. Service has been diverted and restored.
May 24, 2021 11:20 AM EST | Network impact was observed affecting our fibre, DSL and some voice services. Our team is actively investigating and we will post an update as soon as possible
Feb 05, 2021 12:06 AM EST | Splicing complete and all circuits have been reconstructed. Service restored.
Feb 04, 2021 7:25 PM EST | Fibre field techs have identified the failed fibres and available fibres, contractors are on site. They will start pulling cable within the hour. It will probably be a few hours before we are ready to start splicing new cable to restore circuits
Feb 04, 2021 1:00 PM EST | Root cause identified as fibre cut and crews en route to repair the damaged fibre optic cable.
Feb 04, 2021 12:05 PM EST | Issue identified relating to connection to downstream provider. High priority ticket created with cable provider to resolve cross connection issue
Feb 04, 2021 11:40 AM EST | Issue reported for Ontario cable internet customers
Dec 21, 2020 3:30 PM EST | Issue rectified, and we will closely monitor the power consumption to ensure stability of service.
Dec 21, 2020 1:07 PM EST | An issue with inconsistent power at one of our data centres was identified causing intermittent service outages for subset of customers.
Dec 17, 2020 10:00 AM EST | Firewall server issue resolved. All services back to normal
Dec 17, 2020 7:00 AM EST | Firewall server experiencing slowdown. Some customers experiencing slower speeds than usual. Technicians investigating.
Dec 14, 2020 5:45 PM EST | Issue confirmed resolved. If you are still experiencing an issue, please unplug and replug your modem and your connectivity should be restored.
Dec 14, 2020 4:45 PM EST | Hardware failure identified and replacement hardware installed. Service restored.
Dec 14, 2020 2:45 PM EST | Repair ETA 2 hours.
Dec 14, 2020 2:00 PM EST | Technicians are on site and actively working on the issue. We will have an repair ETA shortly.
Dec 14, 2020 1:20 PM EST | Issue identified at one affected data centre – service technicians dispatched
Dec 14, 2020 1:00 PM EST | Regional issue reported, some Internet customers affected
2083 Danforth Ave., 2nd Floor
Toronto, ON, M4C 1K1
Our Contact Information